OpenStudy (anonymous):

Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2.

OpenStudy (anonymous):

y=4x-28 y=4x+28 y=1/4x-2 y=1/4x+2

OpenStudy (anonymous):

I think its B but not 100%

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Here's my reasoning, I know the lines will be parallel because the slope is 4.

OpenStudy (anonymous):

Other than that, it wasa guess. @Loser66

OpenStudy (anonymous):

You reply disappeared on me

OpenStudy (loser66):

It's A , not B. I am sorry, I don't have time to explain

OpenStudy (loser66):

ok, just plug the point into the formula y = mx +b, to find b your x = 8, your y =4 your m = 4 (slope)